OpenStudy (anonymous):

Pre-calculus: At a certain time of day, a 15-meter mast casts a 20 meter shadow. What is the angle of elevation of the sun?

Nnesha (nnesha):



Nnesha (nnesha):

use the trig ratio to find x \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]

OpenStudy (anonymous):

@Nnesha So I'm doing tan because I know that the opposite is 15, and the adjacent is 20? And that's my answer? 15/20?

Nnesha (nnesha):

yes right then use the calculator to find the exact value

OpenStudy (anonymous):

I got .75, is that correct? @Nnesha

Nnesha (nnesha):

remember you should inverse tan \[\huge\rm tan( x) =\frac{ 15 }{ 20 }\] \[x= \tan^{-1} \frac{ 15 }{ 20 }\]

Nnesha (nnesha):

find*

OpenStudy (anonymous):

@Nnesha So the cotangent....? Sorry I just started pre-calc last week so I'm not sure how to do that.

Nnesha (nnesha):

tan and cot are `reciprocal` not inverse use the calculator to find inverse tan

OpenStudy (anonymous):

@Nnesha So tan(15/20) in the calculator?

Nnesha (nnesha):

not tan inverse tan which calculator are u using google or ti84 ?

OpenStudy (anonymous):

@Nnesha Oh I figured it out! The final answer is 36.9. Do you mind explaining why I have to find the inverse?

Nnesha (nnesha):

yes that's right tan and invs are opposite like when you have to find side then you should apply tan and when you have to find `an angle ` then you should use invs tan

Nnesha (nnesha):

i found this from mathisfun website http://prntscr.com/8e6m50

OpenStudy (anonymous):

@Nnesha Thank you SO much. Your help was amazing!

Nnesha (nnesha):

np glad it helped goood work btw! keep it up!